OpenStudy (anonymous):

If someone can scan 1.5 records per minute, how many months does it take them to scan 10,000 records? My friend and I keep getting different answers and our math sequences both look logical.

OpenStudy (anonymous):

\[\frac{(10000/1.5)}{60*24}=4.63 \text{ days} \]

OpenStudy (ash2326):

scanning speed = 1.5 records/minute Total no. of records= 10000 times taken = \(\frac{10000}{1.5} minutes= 6666.67 minutes\) Let's convert minutes into hours => 6666.67/60=111.11 hours let's convert hours into days= 111.11 hours/ 24=4.629 days days into months( assuming 1 month has 30 days)= 4.629/30=0.154 months

OpenStudy (anonymous):

Thank you! It turns out my friend was right, not me.